# {textblock*}

### v1

@@ Abstract
In light of the latest Planck-Atacama Cosmology Telescope (P-ACT) joint results on the primordial scalar power spectrum, we show that the $R^{2}$ inflation model extended with a non-minimally coupled scalar field $\chi$—namely the $R^{2}$-Higgs model—can yield a larger spectral index $n_{s}$ and a small positive running $\alpha_{s}$ at cosmic microwave background (CMB) scales, which are consistent with the data. This is because the $\chi$ field contributes a blue-tilted component to the primordial power spectrum which both modifies the large-scale power and, as a result, significantly enhances power on small scales. The consequent enhancement can be large enough to lead to the formation of primordial black holes (PBHs) of mass $\lesssim 10^{20}\mathrm{g}$ as a dark matter candidate. Furthermore, future observations of the small-scale power spectrum, CMB spectral distortions, and stochastic gravitational waves will provide decisive tests of this model and its predictions for PBHs.
@@ I Introduction
The $R^{2}$ inflation [1, 2], also known as the Starobinsky inflation, has been long regarded as one of the most successful inflationary models, both theoretically well-motivated and perfectly fits the observations. Its predictions for the spectral index $n_{s}$ of the primordial power spectrum and the tensor to scalar ratio $r$ are in excellent agreement with the Planck 2018 [3] and BICEP/Keck observations. The model also gives a precise prediction of a negative running of spectral index $\alpha_{s}$, which is also consistent with the Planck results  [^1] According to the Planck 2018 result, $n_{s}=0.6949\pm 0.0042$ (68% CL),$\alpha_{s}\equiv dn_{s}/d\ln k=-0.0041\pm 0.0067$ (68% CL). The tensor to scalar ratio is constrained to $r<0.032$ (95% CL) by joint Planck and BICEP/Keck observations [88]..
@@
However, a newly released result on the primordial power spectrum from the Atacama Cosmology Telescope (ACT) DR6 [5, 6] reports a slightly larger spectral index. Combined with Planck data at large-to-intermediate scales (denoted as P-ACT) and together with CMB lensing and BAO data from DESI [7] (denoted as P-ACT-LB), the analysis yields a new constraint on the spectral index: $n_{s}=0.9743\pm 0.0034$ (68% CL) at the CMB pivot scale $k_{\mathrm{CMB}}=0.05~{}\mathrm{Mpc}^{-1}$ [5]. This result appears to disfavor the $R^{2}$ model as illustrated in Fig. 4. In addition, the error bar for the running of the spectral index $\alpha_{s}$ is tightened in the P-ACT-LB result, which suggests a slight positive running of the spectral index $\alpha_{s}=0.0062\pm 0.0052$ (68% CL) [6]. The value is not consistent with most of the single-field models that involve a convex potential, including the $R^{2}$ model. Several works [8, 9, 10, 11, 12, 13, 14, 15, 16] discuss extensions to the single-field $R^{2}$ model that may resolve this inconsistency. In this letter, we present a simple extension to the $R^{2}$ model by introducing a non-minimally coupled scalar field $\chi$ (also called $R^{2}$-Higgs model [17, 10, 18, 19, 11]) that may provide a better fit to the newly released data, thus potentially ”save” the $R^{2}$ model via a hybrid-like inflation scenario. The $\chi$ field, acting as a waterfall field, introduces a blue-tilted contribution to the total primordial power spectrum, thereby leading to a positive shift in both the spectral index $n_{s}$ and its running $\alpha_{s}$. Notably, this blue-tilted contribution peaked at small scales can trigger the collapse of overdense regions into primordial black holes (PBHs) [20, 21, 22, 23, 24] during radiation domination, which could constitute cold dark matter.
@@
Therefore, we also discuss the possible hint from the P-ACT results to the small-scale features in the primordial power spectrum. In particular, we find that the scenarios involving the formation of primordial black holes with mass below $10^{20}\mathrm{g}$ through the enhanced small-scale power spectrum, accounting for all of the dark matter, show good agreement with the latest data. The growth of the power spectrum involved in our model also gives rise to distinctive, potentially detectable signatures in both the $\mu$-distortion and the stochastic gravitational wave background, which can serve as distinguishing features in future surveys like PIXIE [25, 26], LISA [27], and etc.
II The $R^{2}$-Higgs model
We consider the action where the $R^{2}$ gravity is non-minimally coupled by a scalar field $\chi$ [17, 10, 18, 19] with a slightly broken $Z_{2}$ symmetry characterized by $\chi_{0}$  [^2] The $Z_{2}$ breaking term is introduced in hybrid models to resolve the overproduction of domain wall and avoid the unacceptably huge quantum fluctuations [89, 19],
@@ II The $R^{2}$-Higgs model
$f(R,\chi)=R+\frac{R^{2}}{6M^{2}}-\frac{\xi R}{M_{\mathrm{pl}}^{2}}(\chi-\chi_{0})^{2},$
@@
$V(\chi)=\frac{1}{4}\lambda(\chi^{2}-m^{2})^{2}.$
Performing conformal transformation $\tilde{g}_{\mu\nu}=Fg_{\mu\nu}$, the action (1) can be transformed into the Einstein frame [30] as
@@
where the $R^{2}$ term converted into a scalar degree of freedom $\phi$, given by
@@
and with an effective potential written as
@@
$W(\phi,\chi)\equiv 1-\frac{1}{F(\phi)}\left[1-\xi\left(\frac{\chi-\chi_{0}}{M_{\mathrm{pl}}}\right)^{2}\right].$
@@
In this framework, a hybrid-like inflation could be realized in the composition of two separated stages of effective single-field inflation dominated individually by $\phi$ and $\chi$ and an intermediate transition stage, which are described as follows:
@@
  • St-1: Effective $R^{2}$ inflation. Initially, the inflaton slow-rolls from the plateau of the potential along the $\phi$ direction. Given the large value of $\phi/M_{\mathrm{pl}}$, the spectator field $\chi$ acquires an effective mass approximately given by $m_{\chi}^{2}=3M^{2}\xi+\mathcal{O}(F^{-1})$.
@@
  • Transition. Effective single-field $R^{2}$ inflation ends when the inflaton reaches the potential minimum in the $\phi$ direction. Simply put, the universe experienced a short but complete version of $R^{2}$ inflation before this moment of transition. During this phase, $F$ approaches unity, $\phi$ becomes heavy ($m_{\phi}^{2}>9H^{2}/4$), and the inflaton dynamics begin to shift toward the $\chi$ direction.
  • St-2: Effective Higgs inflation. The slow roll along $\chi$ starts, with the heavy field $\phi$ stabilized in its potential minimum. The inflation will be ended in effective single-field configuration.
@@
As we explained above, the scalaron $\chi$, acting as a spectator field (or equivalently, the isocurvature component) during the effective $R^{2}$ inflation, dominates the inflation when the St-2 starts. Such background evolution, commonly observed in hybrid inflation models [31, 32, 33] with an initially heavy isocurvature field $\chi$, can induce a sharp $k^{3}$ enhancement in the comoving curvature power spectrum and thus lead to PBH production.
@@
A similar scenario arises in the curvaton model [34, 35, 36, 37, 38]. Unlike the waterfall field in hybrid inflation, the curvaton remains effectively massless and nearly static during inflation. It dominates the universe once the Hubble parameter becomes smaller than the curvaton mass after the end of inflation. In such cases, the comoving curvature power spectrum contributed by curvaton field perturbations that decays later should be close to scale-invariant, depending on the curvaton’s effective mass during inflation.
Notably, both cases are encompassed within the rich phenomenology of the $R^{2}$–Higgs inflationary model. During the first stage of inflation, the effective mass of $\chi$ is given by $m_{\chi}^{2}=F\partial^{2}U/\partial\chi^{2}\sim 3M^{2}\xi$. Since the Hubble parameter in this early stage is approximately $\sim M^{2}/4$, the ratio of $m_{\chi}^{2}$ to $H^{2}$ turns out to solely depend on the coupling constant $\xi$. Therefore, in principle, the model allows the power spectrum to grow from scale-invariant ($k^{0}$) to steeply blue-tilted ($k^{3}$) by varying the coupling constant $\xi$ from 00 to $\gtrsim 3/16$ [19]. Consequently, if a significant enhancement in the power spectrum occurs at small scales, a steeper growth helps to localize the feature, keeping it well separated from large-scale modes. On the other hand, when the spectrum grows more gently, the enhancement at small scales may reshaped the large-scale spectrum.
In the following, we present the $\xi$-dependence of the primordial power spectrum from the perspective of $\delta N$ formalism.
III Analytical power spectrum
@@
For simplicity, we assume that the transition occurs instantaneously. In this case, for modes that exit the horizon before the transition, the total $\delta N$ measured at the end of inflation can be written as
@@
Figure 1: A schematic diagram of the simplified trajectory. We assume that the transition stage occurs instantaneously at $n_{*}$, therefore the background evolution consists of two separate phases of effective single-field inflation along $\phi$ and $\chi$ directions, respectively. The fiducial background trajectory is indicated by black solid arrows, while the perturbed trajectory is shown with red solid arrows. For the perturbed trajectory, the boundary condition perturbed by the field perturbations $\delta\phi_{k}$ (dashed blue arrow) and $\delta\chi_{k}$ (dashed purple arrows) in the spatially flat gauge at the horizon exiting stage of mode $k$, denoted by $n_{k}$.
If we conduct the $\delta N$ calculation nonlinearly, we could observe a large deviation of the probability distribution function of $\mathcal{R}$ from Gaussianity. Such non-Gaussian features are potentially detectable by upcoming observations. We leave this for future work.
@@
during the St-1, which equals the maximum $3$ for the heavy $\chi$ case and smaller when decreasing $\xi$.
Figure 2: The $n_{s}-\alpha_{s}$ relation shown in Eq. (8) for different $M_{\mathrm{PBH}}$s together with the joint constrains from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). To plot the figure, we use $\mathcal{P}_{\mathcal{R}}^{\mathrm{max}}/\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=10^{7}$ in order to have $f_{\mathrm{PBH}}=1$. The hollow black dots refer to cases that $a=3$, or equivalently $\xi\geq 3/16$ when the small-scale amplification is well isolated from the large scales.
As a result, for modes that exit the horizon before the transition stage, the power spectrum contains both contributions from the dominating light field $\phi$ as well as the isocurvature field $\chi$. Assuming a constant $H$ and Bunch-Davis vacuum deep inside of the horizon, the power spectrum for modes $k\leq k_{1}$, where $k_{1}$ is the mode that crosses the horizon during the end of the first phase of inflation (which is approximately the peak scale of the power spectrum), reads
@@
Here the $\mathcal{P}_{\mathcal{R}}^{R^{2}}$ refers to the single field $R^{2}$ model power spectrum, and the peak value of the spectrum $\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}$ is a simple function of $\chi_{0}$ and $\lambda$. Accordingly, the spectral index, the running of the spectral index, and the tensor-to-scalar ratio at the CMB scale are given by
@@
where $N_{1}$ is the number of e-folds before the end of the first stage of inflation corresponding to the CMB pivot scale. Since $k_{1}$ and $N_{1}$ correspond to the peak scale of the power spectrum, considering sharp-peak case, they are related to the PBH mass by
@@
Therefore, to have PBH as all of the dark matter in the present universe, i.e. $f_{\mathrm{PBH}}=1$ is not yet constraint from observation, $N_{1}=30\sim 40$ is needed  [^5] See reviews on the observational constraints on PBH as dark matter [45, 47, 76].
When fixing the ratio of the power spectrum value at the peak scale and the CMB scale to be ${\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}/{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}\sim 10^{7}$ (in order to generate a sizable abundance of primordial black holes today, i.e. $f_{\mathrm{PBH}}\sim 1$ [45, 46, 47]), if the coupling constant $\xi$ is larger than $3/16$ ($a=3$), the exponential suppression renders the last term in $n_{s}$ and $\alpha_{s}$ negligible, which means that the small scale can remain localized without affecting the large scale. In this limit, the large-scale power spectrum reduces to that of the single-field $R^{2}$ inflation, with $\alpha_{s}\lesssim-0.0013$, $n_{s}\lesssim 0.95$ for $N_{1}\lesssim 40$. When $e^{-aN_{1}}$ is of order $10^{-7}$ or less, the modification to the large-scale spectrum comes from the small-scale enhancement can be significant.
@@
$n_{s}(0.05\mathrm{Mpc}^{-1})$ $0.975$ $0.973$ $0.973$
$n_{s}(0.002\mathrm{Mpc}^{-1})$ $0.966$ $0.955$ $0.912$
$\alpha_{s}(0.05\mathrm{Mpc}^{-1})$ $0.008$ $0.014$ 0.0550.0550.0550.055666Excluded at more than $2\sigma$ level.
@@
$f_{\mathrm{PBH}}$ $\sim 1$ $\sim 1$ $\ll 1$
Table 1: The benchmark parameters. The primordial power spectrum is normalized by the Planck 2018 result $\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=\mathcal{P}_{\mathcal{R}}(0.05\mathrm{Mpc}^{-1})=2.1\times 10^{-9}$. A joint P-ACT-LB analysis reports a spectral index $n_{s}(0.05{\rm Mpc}^{-1})=0.9743\pm 0.0034$. The total number of efolds $N$ of the inflation in our model can be tuned from $50\sim 60$ without affecting the IR by adjusting $\lambda$, thus not listed in the table.
Figure 3: The primordial power spectrum measured at the end of inflation together with the (future) observational constraints. The black dashed line shows the result for $\xi>3/16$, whose power spectrum converges to the effective single-field $R^{2}$ inflation. The curves are obtained by numerically solving for the primordial power spectrum by the Transport method using the model parameters from Table 1 case 2 except for a varying $\xi$. From the darkest to the lightest solid curves, we take $\xi=0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (case 2), $0.0685$. The constraint plot comes from [48], including CMB constraints (red) from Planck [3], Lyman-$\alpha$ forest constraints [49] (blue), $\mu$-distortion constraints (green) from FIRAS [50], and PIXIE (future) [25, 26], and gravitational wave constraints (purple) from PTAs [51, 52, 53], SKA [54] (future), LISA [27] (future) and BBO [55] (future).

### v2

@@ Abstract
In light of the latest Planck and Atacama Cosmology Telescope (P-ACT) joint results on the primordial scalar power spectrum, we show that the $R^{2}$ inflation model extended with a non-minimally coupled scalar field $\chi$—namely the $\chi$-extended $R^{2}$ inflation model—can naturally accommodate a larger spectral index $n_{s}$ and a small positive running $\alpha_{s}$ at cosmic microwave background (CMB) scales, both of which are consistent with the latest P-ACT constraints. This is because the $\chi$ field contributes a blue-tilted component to the primordial power spectrum, which both modifies the large-scale power and, as a result, significantly enhances power on small scales. The deviation of the $n_{s}$ and $\alpha_{s}$ from the single field $R^{2}$ inflation is related to the non-minimal coupling constant $\xi$. The consequent enhancement in the primordial power spectrum can be large enough to lead to the formation of primordial black holes (PBHs) of mass $\lesssim 10^{20}\mathrm{g}$ as dark matter candidates. Furthermore, future observations of the small-scale power spectrum, CMB spectral distortions, and stochastic gravitational waves will provide decisive tests of this model and its predictions for PBHs. We stress its strong connection to the seesaw mechanism for the generation of the observed small masses.
@@ I Introduction
[…]
@@
However, a newly released result on the primordial power spectrum from the Atacama Cosmology Telescope (ACT) DR6 [5, 6] reported a slightly larger spectral index. Combined with Planck data at large-to-intermediate scales (denoted as P-ACT) and together with CMB lensing and BAO data from DESI [7] (denoted as P-ACT-LB), the analysis yields a new constraint on the spectral index: $n_{s}=0.9743\pm 0.0034$ (68% CL) at the CMB pivot scale $k_{\mathrm{CMB}}=0.05~\mathrm{Mpc}^{-1}$ [5]. This result appears to disfavor the $R^{2}$ inflation as illustrated in Fig. 5. In addition, the error bar for the running of the spectral index $\alpha_{s}$ is tightened in the P-ACT-LB result, which suggests a slight positive running of the spectral index $\alpha_{s}=0.0062\pm 0.0052$ (68% CL) [6]. The constraint on $\alpha_{s}$ is not consistent with most of the single-field models which involve a convex potential, including the $R^{2}$ model. Several works [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] discussed extensions to the single-field $R^{2}$ model that may resolve this inconsistency.
In this paper, we present a simple extension to the $R^{2}$ model by introducing a non-minimally coupled scalar field $\chi$ (this form is also adopted in $R^{2}$-Higgs model [18, 19, 20, 10, 21, 22, 11]) that may provide a better fit to the newly released data, thus potentially ”save” the $R^{2}$ model via a hybrid-like inflationary scenario. The $\chi$ field, acting as a waterfall field, introduces a blue-tilted contribution to the total primordial power spectrum, thereby leading to a positive shift in both the spectral index $n_{s}$ and its running $\alpha_{s}$. Notably, this blue-tilted contribution peaked at small scales can trigger the gravitational collapse of overdense regions into primordial black holes (PBHs) [23, 24, 25, 26, 27] during radiation domination, which could constitute all cold dark matter.
@@
Therefore, we also discuss the possible hint from the P-ACT results to the small-scale features in the primordial power spectrum. In particular, we find that the scenarios involving the formation of primordial black holes with mass below $10^{20}\mathrm{g}$ through the enhanced small-scale power spectrum, accounting for all of the dark matter, show good agreement with the above latest data. The growth of the power spectrum in our model also gives rise to distinctive, potentially detectable signatures in both the $\mu$-distortion and the stochastic gravitational wave background, which can serve as distinguishing features in future surveys like PIXIE [28, 29, 30], LISA [31, 32, 33], etc.
A crucial point in our model is the presence of a new boson field $\chi$ which has an $Z_{2}$ invariant potential where the $\chi$ is odd under the discrete symmetry. We identify it with the $Z_{4}$ symmetry whose breaking generates the Majorana masses $M_{i}$ for the right-handed neutrinos $N_{Ri}~(i=1-3)$  [34]. Here, the $\chi$ and $N_{Ri}$ have the $Z_{4}$ charges +2+2+ 2 and 1-1- 1, respectively. This $Z_{4}$ is free from the Dai-Freed anomaly [35] and hence it is a perfect symmetry [36, 37] in the standard model. The breaking scale of the $Z_{4}$ is $\langle\chi\rangle\simeq 10^{12}$ GeV which gives the Majorana masses $M_{i}\sim g_{i}\langle\chi\rangle$. Those masses are very much consistent with the observed small masses for the neutrinos [38, 39, 40, 41].
II The $\chi$-extended $R^{2}$ model
We consider the action in which $R^{2}$ gravity is non-minimally coupled to a scalar field $\chi$ [20, 10, 21, 22],
@@ II The $\chi$-extended $R^{2}$ model
$f(R,\chi)=R+\frac{R^{2}}{6M^{2}}-\frac{\xi R}{M_{\mathrm{pl}}^{2}}(|\chi|-\chi_{0})^{2},$
@@
$V(\chi)=\frac{1}{4}\lambda\left(\chi^{2}-v^{2}\right)^{2},$
where $v=\pm m/\sqrt{\lambda}$ is the vacuum expectation value of $\chi$, and $m$ is the mass of $\chi$. Notice that the above action is invariant under the $Z_{4}$ symmetry under which the $\chi$ transforms to $-\chi$. The presence of $|\chi|$ is not a serious problem, since we consider it to be the low-energy action in an effective field theory (EFT). In fact, the $|\chi|$ dependent terms often appear at the EFT level when such a boson field couples to non-abelian gauge theories. (See [43] for example.)
It is well-known that the $R^{2}$ term introduces a scalar degree of freedom to the theory. Introducing an auxiliary field $\psi$, the action (1) is equivalent to,
$S_{\psi}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}RF(\psi,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi$ (2)
$-V(\chi)-U(\psi)\bigg{]},$
where
$F(\psi,\chi)=1+\frac{\psi}{6M^{2}}-\frac{\xi}{M_{\mathrm{pl}}^{2}}(|\chi|-\chi_{0})^{2}$ (3)
$U(\psi)=\frac{M_{\mathrm{pl}}^{2}\psi^{2}}{12M^{2}}$ (4)
To bring the above action into the Einstein frame, we perform the conformal transformation $\tilde{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ [44] for the choice of $\Omega^{2}=F$, where the $\Omega^{2}$ is the conformal factor, and a tilde represents quantities in the Einstein frame. Under transformation, the action (1) can be written in the Einstein frame,
@@
where the scalar field $\phi$ is defined by
@@
equivalently, we define $F(\phi)\equiv\exp(\sqrt{2/3}\phi/M_{\mathrm{pl}})$ and with an effective potential written as
@@
$W(\phi,\chi)\equiv 1-\frac{1}{F(\phi)}\left[1-\xi\left(\frac{|\chi|-\chi_{0}}{M_{\mathrm{pl}}}\right)^{2}\right].$
From now on, we work in the Einstein frame and set $\hbar=c=M_{\mathrm{pl}}=1$, adopting natural units for notational simplicity.
@@
In this framework, a hybrid-like inflation can be realized in the composition of two separated stages of effective single-field inflation — one dominated by $\phi$, the other by $\chi$ — separated by an intermediate transition stage, as described below:
Figure 1: The shape of potential and the background solution with different initial conditions. The solid red and blue lines show the background trajectory obtained by numerically solving the background equations of motion with different initial conditions $\chi_{\mathrm{ini}}>0$ (red) and $\chi_{\mathrm{ini}}<-0$ (blue).
@@
  • St-1: Effective single field $R^{2}$ inflation. Initially, the inflaton slow-rolls from the plateau of the potential along the $\phi$ direction. In the large $\phi$ limit (equivalent to the large $F(\phi)$ limit), the potential becomes
    $U(\phi,\chi)=$ $\frac{3M^{2}}{4}$ (6)
    $+$ $F(\phi)^{-1}\frac{3M^{2}}{2}\left(-1+\xi(|\chi|-\chi_{0})^{2}\right)$
    $+$ $\mathcal{O}(F(\phi)^{-2})V(\chi)$
    Given the large initial value of $\phi$, the spectator field $\chi$ acquires an effective mass approximately given by $m_{\chi}^{2}\approx 3M^{2}\xi(1+\mathcal{O}(F(\phi)^{-1}))$. We assume that the initial value of $\chi$ is close to its potential minimum ±$\pm\chi_{0}$ at the beginning of inflation. During this stage, when $m_{\chi}^{2}>9H^{2}/4\approx 9M^{2}/16$, or equivalently $\xi>3/16$, $\chi$ is considered a heavy field and behaves as an underdamped oscillator. In contrast, when $m_{\chi}^{2}<9H^{2}/4$, $\chi$ is light and behaves as an overdamped oscillator; note that in this case, $\phi$ remains the lightest field. In each of the cases, the scalar field $\chi$ is settled down to one of its potential minimum ($\chi_{0}$ or $-\chi_{0}$, $\chi_{0}$ is taken for the computations in the following session.) before this stage of inflation reaches an end.
@@
  • Transition. The first stage of inflation ends when the inflaton $\phi$ reaches the local potential minimum in the $\phi$ direction. Simply put, the universe undergoes a short but complete version of effective single field $R^{2}$ inflation before this moment of transition. During this phase, $F$ approaches unity when $\phi$ reaches its VEV. The potential at this stage is given in the limit of $|\chi|\rightarrow\chi_{0}$, $\phi\rightarrow 0$ and $|\chi_{0}|\ll m/\sqrt{\lambda}$, the potential is therefore approximately given by:
    $U(\phi,\chi)\approx\frac{m^{4}}{4\lambda}-\frac{1}{2}m^{2}\chi^{2}$ (7)
    As shown above, the boson field $\chi$ becomes unstable $(m_{\chi}^{2}\approx-m^{2}<0)$. Therefore, the inflaton dynamics begin to turn toward the ±$\pm\chi$ direction. It is worth mentioning that $\chi$ field starting from a initial condition away from 0 steps down the potential during the first stage of the $R^{2}$ inflation and hence the $Z_{4}$ symmetry is already broken for the background trajectory, or more precisely,inflation selects a vacuum branch during the early stage (see in Fig. 1). When fixing $|\chi_{0}|^{2}\gg\langle\delta\chi^{2}\rangle$  [45, 46], the overproduction of domain walls (DWs) can be avoided, which usually happens in hybrid inflation models with a discrete $Z_{2}$ symmetry.
  • St-2: Effective single field $\chi$ inflation. The second stage of inflation starts along $\chi$ direction from ±$\pm\chi_{0}$ (the sign will depend on the initial condition for $\chi$), with the heavy field $\phi$ stabilized in its potential minimum. We may observe that $\phi$ is still evolving during this stage, but it is simply due to the varying $\chi$ causing the movement of the potential minimum. The time evolution of $\phi$ is not violating the effective single field slow condition, since the isocurvature is heavy and decays out very rapidly. The second stage of inflation ends as an effective single field inflation of $\chi$. To have a sizable number of efolds during St-2, $v\gg\chi_{0}$ is required. Combining with the non-DW condition, the constraints for the parameters is given
    $m^{2}/\lambda\gg\chi_{0}^{2}\gg M^{2}/(64\pi^{2}).$ (8)
@@
As we explained above, the scalar field $\chi$, acting as a spectator field (or equivalently, the isocurvature component) during the $R^{2}$ inflation, dominates the inflation during St-2. Such background evolution, commonly observed in hybrid inflation models [47, 48, 49] with an initially heavy isocurvature field $\chi$, can induce an enhancement in the comoving curvature power spectrum steepest to $k^{3}$ and thus lead to PBH and GW production.
@@
A similar scenario arises in the curvaton model [50, 51, 52, 53, 54]. In contrast to the waterfall field in hybrid inflation, the curvaton remains effectively massless and nearly frozen during inflation. It dominates the universe once the Hubble parameter becomes smaller than the curvaton mass after the end of inflation. In such cases, the comoving curvature power spectrum contributed by curvaton field perturbations that decays later should be close to scale-invariant, depending on the curvaton’s effective mass during inflation.
Notably, both mechanisms are encompassed within the rich phenomenology of the $\chi$-extended $R^{2}$ model. During the first stage of inflation, the effective mass of $\chi$ is given by $m_{\chi}^{2}=F\partial^{2}U/\partial\chi^{2}\sim 3M^{2}\xi$. Since the Hubble parameter in this early stage is approximately $\sim M^{2}/4$, the ratio of $m_{\chi}^{2}$ to $H^{2}$ turns out to solely depend on the coupling constant $\xi$.
Thus, by varying the coupling constant $\xi$ from 0 to values 3/16$\geq$ 3/16≥ 3 / 16, the model allows the contribution from $\chi$ to the adiabatic power spectrum to evolve from nearly scale-invariant ($k^{0}$) to steeply blue-tilted ($k^{3}$), as explained detailedly in the following section (also shown in [22]).
Consequently, if a significant enhancement in the power spectrum occurs at small scales, a steeper growth helps to localize the feature, keeping it well separated from large-scale modes. On the other hand, when the spectrum grows more gently, the enhancement at small scales may reshape the large-scale spectrum. In the following, we present the $\xi$-dependence of the primordial power spectrum using $\delta N$ formalism.
III Semi-analytical power spectrum
@@
For simplicity, we assume the transition between the two inflationary phases occurs instantaneously.
In this case, for modes that exit the horizon before the transition, the total $\delta N$ evaluated at the end of inflation can be written as
@@
Figure 2: A schematic diagram of the simplified trajectory. We assume that the transition stage occurs instantaneously at $n_{*}$, therefore the background evolution consists of two separate phases of effective single-field inflation along $\phi$ and $\chi$ directions, respectively. The fiducial background trajectory is indicated by black solid arrows, while the perturbed trajectory is shown with red solid arrows. For the perturbed trajectory, the boundary condition is perturbed by the field perturbations $\delta\phi_{k}$ (dashed blue arrow) and $\delta\chi_{k}$ (dashed purple arrows) in the spatially flat gauge at the horizon exiting stage of mode $k$, denoted by $n_{k}$.
If we conduct the $\delta N$ calculation nonlinearly, one may observe a large deviation of the probability distribution function of $\mathcal{R}$ from Gaussianity. Such non-Gaussian features are potentially detectable by upcoming observations. We leave this for future work.
@@
during the St-1, which equals the maximum 333 for the heavy $\chi$ case and smaller when decreasing $\xi$.
Figure 3: The $n_{s}-\alpha_{s}$ relation shown in Eq. (14) for different $M_{\mathrm{PBH}}$s together with the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). To plot the figure, we use $\mathcal{P}_{\mathcal{R}}^{\mathrm{max}}/\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=10^{7}$ in order to have $f_{\mathrm{PBH}}=1$. The hollow black dots refer to cases that $a=3$, or equivalently $\xi\geq 3/16$ when the small-scale amplification is well isolated from the large scales.
As a result, for modes that exit the horizon before the transition stage, the power spectrum contains both contributions from the dominating light field $\phi$ as well as the isocurvature field $\chi$. Assuming a constant $H$ and Bunch-Davis vacuum deep inside the horizon, the power spectrum for modes $k\leq k_{1}$, where $k_{1}$ is the mode that crosses the horizon during the end of the first phase of inflation (which is approximately the peak scale of the power spectrum), is given
@@
Here the $\mathcal{P}_{\mathcal{R}}^{R^{2}}$ refers to the single field $R^{2}$ model power spectrum, and the peak value of the spectrum $\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}$ can be approximate by a simple function of $\chi_{0}$, $\lambda$ and $m$.
Accordingly, the spectral index, the running of the spectral index, and the tensor-to-scalar ratio at the CMB scale are approximately given by
@@
where $N_{1}$ is the number of e-folds before the end of the first stage of inflation corresponding to the CMB pivot scale. Since $k_{1}$ and $N_{1}$ correspond to the peak scale of the power spectrum, considering the sharp-peak case, they are related to the PBH mass by
@@
Therefore, to have PBH as all of the dark matter in the present universe, i.e., $f_{\mathrm{PBH}}=1$ is not yet constraint from observation, $N_{1}=30\sim 40$ is needed  [^4] See reviews on the observational constraints on PBH as dark matter [61, 63, 92].
When the ratio of the power spectrum value at the peak scale and the CMB scale to be ${\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}/{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}\sim 10^{7}$ is fixed to generate a sizable abundance of primordial black holes today, i.e. $f_{\mathrm{PBH}}\sim 1$ [61, 62, 63], if the coupling constant $\xi$ is larger than 3/163/163 / 16 ($a=3$), the modifications from the blue-tilted component to the $n_{s}$ and $\alpha_{s}$ at CMB scale is exponentially suppressed, effectively decoupling small-scale features from large-scale observables. In this limit, the large-scale power spectrum reduces to that of the single-field $R^{2}$ inflation, with $\alpha_{s}\lesssim-0.0013$, $n_{s}\lesssim 0.95$ for $N_{1}\lesssim 40$. When $e^{-aN_{1}}$ is of order $10^{-7}$ or less, the modification to the large-scale spectrum comes from the small-scale enhancement can be significant.
@@
$n_{s}(0.05\mathrm{Mpc}^{-1})$ 0.9750.9750.975 0.9730.9730.973 0.9730.9730.973
$n_{s}(0.002\mathrm{Mpc}^{-1})$ 0.9660.9660.966 0.9550.9550.955 0.9120.9120.912
$\alpha_{s}(0.05\mathrm{Mpc}^{-1})$ 0.0080.0080.008 0.0140.0140.014 0.0550.0550.055555Excluded at more than $2\sigma$ level.
@@
$f_{\mathrm{PBH}}$ 1$\sim$ 1∼ 1 1$\sim$ 1∼ 1 1$\ll$ 1≪ 1
Table 1: The benchmark parameters and the observables. The primordial power spectrum is normalized by the Planck 2018 result $\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=\mathcal{P}_{\mathcal{R}}(0.05\mathrm{Mpc}^{-1})=2.1\times 10^{-9}$. A joint P-ACT-LB analysis reported a spectral index $n_{s}(0.05{\rm Mpc}^{-1})=0.9743\pm 0.0034$. The total number of efolds $N$ of the inflation in our model can be tuned from 506050$\sim$ 6050 ∼ 60 without affecting the IR by adjusting $\lambda$, thus not listed in the table.
[…]